OpenStudy (anonymous):

The system, 6x - 4y = 5 and 9x - 6y = 5, is consistent. Answer True False

OpenStudy (anonymous):

false

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

because when we solve them by putting x=(5+4y)/6 in second equation we get 9((5+4y)/6-6y=5 or 45+36y-36y=5 or 45=5 which is false

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

2x + y = 7 and x - 2y = 7 so this would be consistent